A 50 kg child stands at the rim of a merry-go-round with radius 2 meters, rotating with an angular speed of 3 rad/s.

(a) What is the child's tangential (linear) speed? Answer - 6 m/s (show how to get)

(b) What is the child's centripetal acceleration? Answer 18 m/s2 (show how to get)

(c) What provides the force between her feet and floor? Answer Friction. (explain)

(d.) What is the force between her feet and the floor carousel? answer 900N (show how to get)
Answer #1

a) The tangential speed, v = ω*r

Given, ω = 3 rad/s, r = 2 m

So, v = 3 * 2 = 6 m/s (Answer)

b) The centripetal acceleration is given by,

                a = v²/r

i.e. a = (6)²/2 = 36/2 = 18 m/s² (Answer)